OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 8x + 5y = 3 7x – 3y = 10

jhonyy9 (jhonyy9):

multiply first by 3 and second by 5 24x+15y=9 35x-15y=50 ------------ 59x 0 =59 x=1 24*1 +15y=9 15y=9-24 15y=-15 y=-15/15 y=-1 x=1 y=-1